# TIDMAD: Time Series Dataset for Discovering Dark Matter with AI Denoising

### v2

@@ 5 Limitations and applications
As illustrated in Figure 4.2, the baseline models fail to surpass the results of ABRA-10cm Run 3. This is attributed to hardware limitations and changes since the last data run, including (1) replacing the dark matter pickup cylinder with a pickup loop, consequently reducing the geometric coupling to the dark matter signal, and (2) reducing the data taking period to 24 hours from three months. The decision to implement these changes was driven by the aim to enable ABRA-10cm to simultaneously search for dark matter and gravitational waves, thereby enhancing the scientific scope of the experiment. The shortened data taking duration was necessitated by operational constraints of the dilution refrigerator. Because the signal-to-noise ratio scales as the fourth root of integration time, we can increase ABRA-10cm’s sensitivity to dark matter by increasing the data taking period [18]. Another more efficacious way to increase this signal-to-noise ratio is improving our denoising with ML; a doubling of our noise reduction represents a 16x speed up our data taking time revealing the out-sized return on investment in denoising techniques when compared to increased detector run time.
@@
In Section 4.2, we discussed how to set a dark matter limit using the provided analysis scripts. This script assumes a null result as no $5\sigma$ dark matter candidates were identified in this region of parameter space by ABRA-10cm Run 3. Therefore, we assume a null result for this much shorter ($24$ hr) data taking. This assumption enables us to establish upper limits on the coupling parameter $g_{a\gamma\gamma}$ for every mass point. As shown in Figure 4.2, the ABRA-TIDMAD Raw limit without denoising covers a smaller region than ABRA-10cm Run 3 [25].
However, with the denoising algorithm applied to the 24-hour science data, ABRA-TIDMAD could potentially reach beyond the ABRA-10cm Run 3 region, where a discovery of dark matter is possible. In this paper, we focused on increasing the experimental sensitivity and setting exclusion limits. A straightforward modification to the interpretation of the TS would unlock the discovery potential of this analysis framework. Future efforts will focus on employing a more extensive dataset and implementing dark matter discovery analysis code. Given that discovering dark matter would be a Nobel Prize-level breakthrough, it is crucial to not only claim discovery but also to convince the scientific community of its validity. If TIDMAD users find any anomalous signals in the science dataset, please contact the authors for further investigation and understandings of systematic uncertainties.
@@
Axion dark matter is a specific subset of wave-like dark matter candidates, making the techniques developed in this paper broadly applicable to a wide range of wave-like dark matter experiments. Other axion dark matter experiments, including but not limited to ADMX [7], HAYSTAC [6], and CASPEr [8], also produces long time series data and search for similar peaks in frequency domain; any AI algorithm developed upon TIDMAD can be easily adapted and applied to these experiments. Furthermore, time series denoising algorithms are crucial for extracting wave-like signals in various areas of physics. In astrophysics, gravitational wave searches involve detecting chirp signals with durations on the order of seconds [1], often buried within detector noise. In nuclear physics, denoising can enhance the efficiency of HPGe detectors [5] and bolometer detectors [32]. Advancements in denoising can be deployed across a suite of these physics experiments.
@@ 6 Conclusions and other works
We present TIDMAD, the first dataset and benchmark designed to yield a community-standard dark matter search result. TIDMAD includes all necessary inputs and processing to train time series denoising algorithms and produce a science-level dark matter limit. Through a series of experiments, we developed three ultra-long time series deep learning algorithms, benchmarked their ability to recover hardware-injected signals, and set dark matter limits. Clear performance improvements were demonstrated on both benchmarks. Our future work will focus on enhancing the denoising algorithm to achieve better dark matter limits, expanding to other nuclear and particle experiments, and embedding these algorithms onto FPGA chips for real-time denoising during data taking.
@@ Appendix A Dark Matter Signal and Signal Injection
Axion dark matter appears in our detector as a time oscillating current with two free model parameters ($m_{a},g_{a\gamma\gamma}$) described by Equation 1. This signal can be fully derived from axion theory and is extremely well-defined. While a full derivation of this equation is presented in Reference [18], schematically, this theoretical prediction for the dark matter signal in our detection comes from four phenomena -- the omnipresent axion field, axion’s interactions with electromagnetism, the specific geometry of our detector, and the velocity distribution of the dark matter.
@@
The axion field can interact with other physical forces including electromagnetism. If an axion field collides with an electromagnetic field, some of the axions will be transformed into photons. These photons produce a secondary electromagnetic field -- an axion-induced electromagnetic field. In the geometry of our detector, the axion-induced electromagnetic field is read out with a pickup wire which samples the field like a typical radio antenna. This means the time oscillating axion field turns into a time oscillating current in our detector (i.e. our dark matter signal).
@@
The second free signal parameter is the sine wave amplitude which is proportional to $g_{a\gamma\gamma}$, the strength of the axion’s interactions with electromagnetism. Theoretical calculations constrain this parameter to $g_{a\gamma\gamma}=Cm_{a}$ where C=[0.39,0.22]C=[-0.39,0.22] depending on the theory [26, 13]. Theoretically motivated axion couplings can be seen in Figure 4.2 as the gold band in the $m_{a},g_{a\gamma\gamma}$ parameter space. Experimentally, the goal is to detect ever smaller signal amplitudes to reach lower values of $g_{a\gamma\gamma}$.
@@
To summarize, our signal injection scheme involves exciting hardware with a sine wave from a signal generator. We sequentially step through 309 different frequencies from 1.1 kHz to 4.9 MHz to simulated 309 different axion masses in our detector hardware all with an amplitude of 50 mV so that the fake signal is visible above the detector noise floor.
@@ Appendix B Details and limitations of deep learning model
There are two special treatments we took to train the three deep learning models:
@@
Since the injected dark matter signal spans two orders of magnitude in frequency, the observed features in the injected time series significantly vary. During training, we noticed that a single deep learning model to denoise the entire dataset would fail to generalize across the different injected frequency ranges. To address this issue, we trained four deep learning models per architecture, each focusing on a specific frequency range: the first covering the low-frequency regime (training/validation files 0-3), the second covering the mid-low regime (training/validation files 4-9), the third covering the mid-high regime (training/validation files 10-14), and the fourth covering the high-frequency regime (training/validation files 15-19). During the benchmark 1 inference, we selected the input validation data corresponding to the frequency range for which each model was trained and averaged the results of the four models. For the benchmark 2 inference, we ran all four models on the science data and selected the highest-performing model, as represented in Figure 4.2. In future work, we aim to develop a single denoising model generalizable to wide frequency ranges. All trained models are made available in https://drive.google.com/drive/folders/16ORX1b2zo1_lOYYAcRBgddBuYImj0Bxs?usp=share_link.
The hyperparameter of FC Net is listed below:
@@
The output of FC Net at every time step is a single float point number. An MSE loss is calculated between the float point number and the ground truth value.
The PU Net model consists of four down layers and four up layers, with contracting paths between each pair of layers. The down layers include Max Pooling and two convolutional operations, while the up layers comprise Deconvolution and Convolution. Additionally, positional encoding is added after each down layer  [31]. Lastly, the output is fed into a linear layer to produce $256$-dimensional vector at each time step The detailed model hyperparameter could be found within the network.py script in https://github.com/jessicafry/TIDMAD.
@@
The WaveNet model implements the core dilated causal convolution architecture with 10 residual blocks using exponentially increasing dilation rates (1, 2, 4, 8, 16, ..., 512), kernel size of 12, and 32 residual channels with 64 gate channels for the gated activation mechanism (tanh × sigmoid). Each block processes input through dilated causal convolutions to ensure no future information leakage, applies gated activation, and produces both residual connections (for depth) and skip connections (accumulated across all blocks for the final output). The model uses 32 skip channels and projects the final accumulated skip connections through two 1D convolutions to produce the output shape of (batch, sequence_length, 256) while achieving a large receptive field of over 1000 timesteps through the dilated convolution hierarchy, making it suitable for modeling long-term dependencies in sequential data like audio waveforms.
@@ Appendix C Frequentist log-likelihood test statistics
The detailed analysis flow to produce the dark matter limit is depicted in Figure 7. The first step involves performing a fast Fourier transform on the time series data in 10-second segments to produce power spectral densities (PSDs). These PSDs are then averaged across the full dataset to generate our average PSD, reflecting the power in the pickup loop as a function of frequency. Since one of the physics parameters, dark matter mass ($m_{a}$), , is directly proportional to the frequency, the analysis script conducts 11.1 million independent searches for dark matter with varying mass points from 0.4 neV (100 kHz) to 8 neV (2 MHz).
@@ Appendix D Current Axion Limits
In Figure 4.2, the TIDMAD dark matter limits, denoised and raw, are presented alongside ABRA-10cm’s previous Run 3 limits and various state-of-the-art axion dark matter experiments and observations. In gold, the theoretically motivated couplings for dark matter candidates are highlighted as discussed in Appendix A. In this appendix, a brief summary of the origins of these limits will be provided.
@@
In this class of experiments, a high-intensity laser is directed towards a solid barrier. While conventional light cannot traverse the barrier, the interaction with the wall may cause a fraction of the light to convert into axions. As a result of their weak interaction with matter, these axions could pass through the wall. On the opposite side, detectors are placed to identify any reconverted light, which would indicate the presence of axions. Light-shining-through-walls experiments must both create axions from photons and detect these axions by converting them back into photons, whereas ABRA-10cm only needs to detect axions, not create them.
@@
This class of experiments generate a strong magnetic field to stimulate axions to convert into microwave photons within a resonant cavity-enclosed space. The resonant cavity is finely tuned to amplify specific frequencies of electromagnetic radiation. Sensitive radio receivers then measure the power within the cavity to identify any potential photon signals indicative of axions. While the conversion mechanism is identical to ABRA-10cm, the resonant cavity only amplifies targeted frequencies, while ABRA-10cm’s readout chain has broadband amplification of axion induced signals. ABRA-10cm, Cavity Haloscopes, and SHAFT are all examples of Haloscopes -- experiments that search for dark matter axions in our galaxy’s dark matter halo.
@@ E.2 Result and Analysis
The results from the 10 Moving Average trials are presented in Table 2. The complete results from all 50 Savitzky–Golay filter trials are displayed in Figure 4.2, with a representative subset also included in Table 2.
@@
Savitzky–Golay Filter Performance: The algorithm achieves better performance with smaller window sizes and lower polynomial orders. Notably, the best denoising score was achieved in a trial with window size 19 and polynomial order 11. This means the SG fitter fits an 11-dimensional polynomial to every 19 elements in the time series --- representing an extreme overfitting scenario. In other words, SG filter in this trial attempts to memorize and preserve all features in the time series, resulting in a Fine Denoising Score of 0.94, which asymptotically approach but still inferior to the baseline score of 1.0 (when no denoising algorithm is applied).
@@ E.3 Conclusion and Key Finding
NeurIPS Paper Checklist
@@
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the claims made in the paper. The claims in the introduction, abstract, and main paper are the same; we introduce a new, segmented dataset from a physics detector, a denoising score for direct model benchmarking, and a complete dark matter analysis framework for science legible benchmarking.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
@@ 5. Guidelines:
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Included in Section 5 is a discussion of the limitation of this work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
@@ 10. Guidelines:
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section 1 and Section A, we discuss the physics theory and assumptions for our result. Additionally, for the evaluation metric, we provide a set of assumptions for the benchmark construction in Section 4.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.

### v3

@@ 5 Limitations and applications
As illustrated in Figure 5, the baseline models fail to surpass the results of ABRA-10cm Run 3. This is attributed to hardware limitations and changes since the last data run, including (1) replacing the dark matter pickup cylinder with a pickup loop, consequently reducing the geometric coupling to the dark matter signal, and (2) reducing the data taking period to 24 hours from three months. The decision to implement these changes was driven by the aim to enable ABRA-10cm to simultaneously search for dark matter and gravitational waves, thereby enhancing the scientific scope of the experiment. The shortened data taking duration was necessitated by operational constraints of the dilution refrigerator. Because the signal-to-noise ratio scales as the fourth root of integration time, we can increase ABRA-10cm’s sensitivity to dark matter by increasing the data taking period [20]. Another more efficacious way to increase this signal-to-noise ratio is improving our denoising with ML; a doubling of our noise reduction represents a 16x speed up our data taking time revealing the out-sized return on investment in denoising techniques when compared to increased detector run time.
@@
In Section 4.2, we discussed how to set a dark matter limit using the provided analysis scripts. This script assumes a null result as no $5\sigma$ dark matter candidates were identified in this region of parameter space by ABRA-10cm Run 3. Therefore, we assume a null result for this much shorter ($24$ hr) data taking. This assumption enables us to establish upper limits on the coupling parameter $g_{a\gamma\gamma}$ for every mass point. As shown in Figure 5, the ABRA-TIDMAD Raw limit without denoising covers a smaller region than ABRA-10cm Run 3 [28].
However, with the denoising algorithm applied to the 24-hour science data, ABRA-TIDMAD could potentially reach beyond the ABRA-10cm Run 3 region, where a discovery of dark matter is possible. In this paper, we focused on increasing the experimental sensitivity and setting exclusion limits. A straightforward modification to the interpretation of the TS would unlock the discovery potential of this analysis framework. Future efforts will focus on employing a more extensive dataset and implementing dark matter discovery analysis code. Given that discovering dark matter would be a Nobel Prize-level breakthrough, it is crucial to not only claim discovery but also to convince the scientific community of its validity. If TIDMAD users find any anomalous signals in the science dataset, please contact the authors for further investigation and understanding of systematic uncertainties.
@@
Axion dark matter is a specific subset of wave-like dark matter candidates, making the techniques developed in this paper broadly applicable to a wide range of wave-like dark matter experiments. Other axion dark matter experiments, including but not limited to ADMX [8], HAYSTAC [7], and DMRadio [9], also produces long time series data and search for similar peaks in frequency domain; any AI algorithm developed upon TIDMAD can be easily adapted and applied to these experiments. Because the denoising and benchmarking methods are detector-agnostic, the authors are presently collaborating with several other axion dark matter experiments beyond ABRA-10cm on implementing the denoising procedure presented in this paper to various experimental set ups. This work lays the groundwork for such generalization, and we hope the paper will encourage similar releases from other experimental collaborations. Furthermore, time series denoising algorithms are crucial for extracting wave-like signals in various areas of physics. In astrophysics, gravitational wave searches involve detecting chirp signals with durations on the order of seconds [1, 23], often buried within detector noise. In nuclear physics, denoising can enhance the efficiency of HPGe detectors [5] and bolometer detectors [35]. Advancements in denoising can be deployed across a suite of these physics experiments.
@@ 6 Conclusions and other works
We present TIDMAD, the first dataset and benchmark designed to yield a community-standard dark matter search result. TIDMAD includes all necessary inputs and processing to train time series denoising algorithms and produce a science-level dark matter limit. Through a series of experiments, we developed five ultra-long time series deep learning algorithms, benchmarked their ability to recover hardware-injected signals, and set dark matter limits. Clear performance improvements were demonstrated on both benchmarks. Our future work will focus on enhancing the denoising algorithm to achieve better dark matter limits, expanding to other nuclear and particle experiments, and embedding these algorithms onto FPGA chips for real-time denoising during data taking.
@@ Appendix A Dark Matter Signal and Signal Injection
Axion dark matter appears in our detector as a time oscillating current with two free model parameters ($m_{a},g_{a\gamma\gamma}$) described by Equation 1. This signal can be fully derived from axion theory and is extremely well-defined. While a full derivation of this equation is presented in Reference [20], schematically, this theoretical prediction for the dark matter signal in our detection comes from four phenomena – the omnipresent axion field, axion’s interactions with electromagnetism, the specific geometry of our detector, and the velocity distribution of the dark matter.
@@
The axion field can interact with other physical forces including electromagnetism. If an axion field collides with an electromagnetic field, some of the axions will be transformed into photons. These photons produce a secondary electromagnetic field – an axion-induced electromagnetic field. In the geometry of our detector, the axion-induced electromagnetic field is read out with a pickup wire which samples the field like a typical radio antenna. This means the time oscillating axion field turns into a time oscillating current in our detector (i.e. our dark matter signal).
@@
The second free signal parameter is the sine wave amplitude which is proportional to $g_{a\gamma\gamma}$, the strength of the axion’s interactions with electromagnetism. Theoretical calculations constrain this parameter to $g_{a\gamma\gamma}=Cm_{a}$ where C=[0.39,0.22]C=[-0.39,0.22] depending on the theory [29, 14]. Theoretically motivated axion couplings can be seen in Figure 5 as the gold band in the $m_{a},g_{a\gamma\gamma}$ parameter space. Experimentally, the goal is to detect ever smaller signal amplitudes to reach lower values of $g_{a\gamma\gamma}$.
@@
To summarize, our signal injection scheme involves exciting hardware with a sine wave from a signal generator. We sequentially step through 309 different frequencies from 1.1 kHz to 4.9 MHz to simulate 309 different axion masses in our detector hardware all with an amplitude of 50 mV so that the fake signal is visible above the detector noise floor.
@@ Appendix B Details and limitations of deep learning model
There are two special treatments we took to train the five deep learning models:
@@
Since the injected dark matter signal spans two orders of magnitude in frequency, the observed features in the injected time series significantly vary. During training, we noticed that a single deep learning model (with the exception of WaveNet) to denoise the entire dataset would fail to generalize across the different injected frequency ranges. To address this issue, we trained four deep learning models per architecture, each focusing on a specific frequency range: the first covering the low-frequency regime (training/validation files 0-3), the second covering the mid-low regime (training/validation files 4-9), the third covering the mid-high regime (training/validation files 10-14), and the fourth covering the high-frequency regime (training/validation files 15-19). During the benchmark 1 inference, we selected the input validation data corresponding to the frequency range for which each model was trained and averaged the results of the four models. For the benchmark 2 inference, we ran all four models on the science data and selected the highest-performing model, as represented in Figure 5.
The WaveNet architecture was the only deep learning model that generalized across the entire frequency range of validation data. All trained models are made available in https://drive.google.com/drive/folders/16ORX1b2zo1_lOYYAcRBgddBuYImj0Bxs?usp=share_link.
The hyperparameters of the FC Net are listed below:
@@
The output of FC Net at every time step is a single floating point number. An MSE loss is calculated between the floating point number and the ground truth value.
The PU Net model consists of four down layers and four up layers, with contracting paths between each pair of layers. The down layers include Max Pooling and two convolutional operations, while the up layers comprise Deconvolution and Convolution. Additionally, positional encoding is added after each down layer  [34]. Lastly, the output is fed into a linear layer to produce $256$-dimensional vector at each time step. The detailed model hyperparameter could be found within the network.py script in https://github.com/jessicafry/TIDMAD.
@@
The WaveNet model implements the core dilated causal convolution architecture with 10 residual blocks using exponentially increasing dilation rates (1, 2, 4, 8, 16, …, 512), kernel size of 12, and 32 residual channels with 64 gate channels for the gated activation mechanism (tanh × sigmoid). Each block processes input through dilated causal convolutions to ensure no future information leakage, applies gated activation, and produces both residual connections (for depth) and skip connections (accumulated across all blocks for the final output). The model uses 32 skip channels and projects the final accumulated skip connections through two 1D convolutions to produce the output shape of (batch, sequence_length, 256) while achieving a large receptive field of over 1000 timesteps through the dilated convolution hierarchy, making it suitable for modeling long-term dependencies in sequential data like audio waveforms.
@@ Appendix C Frequentist log-likelihood test statistics
The detailed analysis flow to produce the dark matter limit is depicted in Figure 7. The first step involves performing a fast Fourier transform on the time series data in 10-second segments to produce power spectral densities (PSDs). These PSDs are then averaged across the full dataset to generate our average PSD, reflecting the power in the pickup loop as a function of frequency. Since one of the physics parameters, dark matter mass ($m_{a}$), is directly proportional to the frequency, the analysis script conducts 11.1 million independent searches for dark matter with varying mass points from 0.4 neV (100 kHz) to 8 neV (2 MHz).
@@ Appendix D Current Axion Limits
In Figure 5, the TIDMAD dark matter limits, denoised and raw, are presented alongside ABRA-10cm’s previous Run 3 limits and various state-of-the-art axion dark matter experiments and observations. In gold, the theoretically motivated couplings for dark matter candidates are highlighted as discussed in Appendix A. In this appendix, a brief summary of the origins of these limits will be provided.
@@
In this class of experiments, a high-intensity laser is directed towards a solid barrier in a magnetic field. While conventional light cannot traverse the barrier, the interaction with the magnetic field may cause a fraction of the light to convert into axions. As a result of their weak interaction with matter, these axions could pass through the wall. On the opposite side, detectors are placed to identify any reconverted light, which would indicate the presence of axions. Light-shining-through-walls experiments must both create axions from photons and detect these axions by converting them back into photons, whereas ABRA-10cm only needs to detect axions, not create them.
@@
This class of experiments generate a strong magnetic field to stimulate axions to convert into microwave photons within a resonant cavity-enclosed space. The resonant cavity is finely tuned to amplify specific frequencies of electromagnetic radiation. Sensitive radio receivers then measure the power within the cavity to identify any potential photon signals indicative of axions. While the conversion mechanism is identical to ABRA-10cm, the resonant cavity only amplifies targeted frequencies. In contrast, ABRA-10cm’s readout chain has broadband amplification of axion induced signals. ABRA-10cm, Cavity Haloscopes, and SHAFT are all examples of Haloscopes – experiments that search for dark matter axions in our galaxy’s dark matter halo.
@@ E.2 Result and Analysis
The results from the 10 Moving Average trials are presented in Table 2. The complete results from all 50 Savitzky–Golay filter trials are displayed in Figure 5, with a representative subset also included in Table 2.
@@
Savitzky–Golay Filter Performance: The algorithm achieves better performance with smaller window sizes and lower polynomial orders. Notably, the best denoising score was achieved in a trial with window size 19 and polynomial order 11. This means the SG fitter fits an 11-dimensional polynomial to every 19 elements in the time series — representing an extreme overfitting scenario. In other words, SG filter in this trial attempts to memorize and preserve all features in the time series, resulting in a Fine Denoising Score of 0.94, which asymptotically approach but still inferior to the baseline score of 1.0 (when no denoising algorithm is applied).
@@ E.3 Conclusion and Key Finding
TIDMAD: Supplemental Information
TIDMAD: Supplemental Information
Appendix 1 Datasheet
1.1 Motivation
@@
    For what purpose was the dataset creates? Our datasets were created to train and benchmark ultra-long time series denoising frameworks for the discovery of dark matter.
@@
    Who created the dataset and on behalf of which entity? This dataset is the direct output of the ABRACADABRA detector on behalf of the researchers on the author list. The ABRACADABRA detector was built by the ABRACADABRA collaboration including J. T. Fry and the authors of [28].
@@
    Who funded the creation of the dataset? This work was generously funded by the National Science Foundation under grant numbers NSF-PHY-1658693, NSF-PHY-1806440. J. T. Fry is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064
1.2 Composition
  1. 1.
    What do the instances that comprise the dataset represent? Each instance represents a voltage at a moment in time read out by our detector. For the SQUID data, this voltage comes from flux on the pickup loop of wire, converted to a voltage by the SQUID detector, read out by our digitizer. For the SG data, this voltage comes directly from a signal generator passed through a power splitter.
  2. 2.
    How many instances are there in total? There are 867,260,000,000 voltage instances total.
  3. 3.
    Does the dataset contain all possible instances or is it a sample of instances from a larger set? The voltage produced by the SQUID and the SG are continuous. The instances are sampled from this continuous voltage stream at a constant rate of 10MS/s.